OpenStudy (anonymous):

Does 7+28+112+448+... converge or diverge?

OpenStudy (psymon):

This is a series I can put into the form of ar^n, where a is the initial value and r^n is basically the pattern of the series. I can fully write the series like this: \[\sum_{i=0}^{\infty}ar ^{n}\] Where 7 is your a and 4 is your r, because that is the pattern here, multiplying by 4. \[\sum_{i=0}^{\infty}7(4)^{n}\] Now when the absolute value of r, 4 in this case, is greater than or equal to 1, your series diverges. So this series diverges. If your absolute value of r was 0<r<1, the series would converge. If your series converges, your sum is equal to: \[\frac{ a }{ 1-r }\] Of course it diverges here so you don't need that, but I thought I would put it :P

OpenStudy (anonymous):

thank you

OpenStudy (psymon):

Np :3